# MUTE-SLAM: Real-Time Neural SLAM with Multiple Tri-Plane Hash Representations

### v1

@@ Abstract
We introduce MUTE-SLAM, a real-time neural RGB-D SLAM system employing multiple tri-plane hash-encodings for efficient scene representation. MUTE-SLAM effectively tracks camera positions and incrementally builds a scalable multi-map representation for both small and large indoor environments. It dynamically allocates sub-maps for newly observed local regions, enabling constraint-free mapping without prior scene information. Unlike traditional grid-based methods, we use three orthogonal axis-aligned planes for hash-encoding scene properties, significantly reducing hash collisions and the number of trainable parameters. This hybrid approach not only ensures real-time performance but also enhances the fidelity of surface reconstruction. Furthermore, our optimization strategy concurrently optimizes all sub-maps intersecting with the current camera frustum, ensuring global consistency. Extensive testing on both real-world and synthetic datasets has shown that MUTE-SLAM delivers state-of-the-art surface reconstruction quality and competitive tracking performance across diverse indoor settings. The code will be made public upon acceptance of the paper.
Keywords:
Tri-plane Hash-Encoding Multi-Map Representation Scalability
1 Introduction
@@
With the advent of Neural Radiance Field (NeRF) [16], efforts have been made to integrate it into SLAM systems due to its capability to render novel views [2, 3, 4, 31] and reconstruct 3D surfaces [1, 21, 30, 14]. NeRF-based SLAM methods iMAP [25] and NICE-SLAM [35], have demonstrated their applicability across various scenes and ability to predict the appearance of unobserved areas, although their computational demands hinder real-time application. Recent works [28, 11] utilize hash-encoded voxel grids [17] to accelerate convergence and enhance detail fidelity. Despite hash collisions can be implicitly mitigated by the original design of Instant-NGP [17], the reconstructed mesh still suffers from aliasing. As proved in [36, 12, 15, 13, 15], projecting spatial points into a tri-plane reprensentation can efficiently reduce scene parameters while preserving geometric and color information. Therefore following [13], we leverage a tri-plane hash-encoding method to store scene features and minimize hash collisions. Moreover, current Nerf-based SLAM systems [25, 35, 12, 28] require pre-set scene boundaries, limiting their application in unknown environments. Some [11, 33] address this with octree-based voxel grids, but they still necessitate an initially defined loose boundary and struggle to reconstruct beyond these limits. Our proposed MUTE-SLAM overcomes this by introducing a multi-map-based scene representation. Through dynamically allocating new sub-maps upon detecting new areas, MUTE-SLAM can be deployed in environments of any size given reasonable RGB-D sensor observations.
@@
2 Related Works
2.1 Dense Visual SLAM
@@
2.2 Neural Implicit SLAM
@@
Recent advancements in NeRF-based SLAM have improved tracking, mapping, and running speed. ESLAM [12] stores features on multi-scale axis-aligned planes and employs rendering based on TSDF. Co-SLAM [28] combines coordinate and hash-encodings for input points to achieve both smooth meshes and fast convergence, and introduces a real-time global bundle adjustment mechanism using a ray list sampled from past keyframes. Despite these improvements, scalability remains a challenge. Our work addresses this by introducing a multi-map solution with an accompanying optimization strategy. Besides, we represent each sub-map with a tri-plane hash-encoding which allows for fast convergence and detailed surface reconstruction. Concurrent work [8] also adopts a tri-plane multi-map representation, but differs from ours in using a joint coordinate and tri-plane encoding, as well as in local map allocation and optimization strategies.
@@
3 Method
The overview of MUTE-SLAM is illustrated in Fig. 2. Given an input RGB-D stream $\{I_{i},D_{i}\}_{i=1}^{M}$, the system tracks their 6-DOF camera poses $\{T_{i}:R_{i}|t_{i}\}_{i=1}^{M}$ and conducts mapping to optimize a multi-map implicit scene representation $\{E_{k}\}_{k=1}^{S}$. When the camera captures a new region, a corresponding sub-map is created to represent it, as detailed in Sec. 3.1. Each local map $E_{k}$ encodes a point coordinate $\bm{p}$ within its domin with three orthogonal TSDF feature planes and three color planes, as described in Sec. 3.2. These features from the sub-maps are decoded into TSDF and color using two separate MLPs, initiating the volume rendering process (Sec. 3.3). The rendered depth and color images are subsequently utilized to jointly optimize the sub-maps and camera poses. Additionally, periodic global bundle adjustments are implemented to ensure global consistency(Sec. 3.4).
3.1 Multi-map Scene Representation
@@
3.2 Tri-plane Hash Encoding
@@
The TSDF and color values are then utilized in the volume rendering module. The plane-based representation, growing quadratically with scene size, results in fewer hash table queries and hence fewer collisions compared to grid representations, given equal hash table sizes. Furthermore, since a point’s feature vector is a composite of inputs from three distinct encoders, the probability of encountering conflicts across all encoders is significantly reduced. In situations where collisions do occur in one encoder, the impact is mitigated by the inputs from the other encoders, thus lessening the overall adverse effect. Sec. 4.4 demonstrates the effectiveness of our tri-plane hash-encoding.
3.3 TSDF-based Volume Rendering
@@
[…]
@@
3.4 Tracking and Mapping
3.4.1 Loss Functions.
[…]
@@ 3.4 Tracking and Mapping
[…]
@@
4 Experiments.
4.1 Experimental Setup.
4.1.1 Baselines.
We choose state-of-art NeRF-based dense SLAM approaches ESLAM [12] and Co-SLAM [28] as our main baselines for both surface reconstruction and camera tracking. To better evaluate our proposed MUTE-SLAM on pose estimation, we also compare with previous methods NICE-SLAM [35] and Vox-Fusion [33]. We run these methods using the default settings provided in their open-source code.
4.1.2 Datasets.
@@
4.1.3 Metrics.
For surface reconstruction, We adopt evaluation metrics $Depth\ L1\ [cm]$, $Accuracy\ [cm]$, $Completion\ [cm]$ and $Completion\ ratio\ [<5cm\%]$. Additionally, to underscore the ability of our method to produce detailed geometry compared to ESLAM [12], we incorporate the $Completion\ ratio\ [<1cm\%]$ metric. Following [1, 29, 12], before evaluation, we remove faces that are not inside any camera frustum or are occluded in all RGB-D frames from the reconstructed mesh. For the evaluation of camera tracking, we employ $ATE\ RMSE\ [cm]$ [24] .
4.1.4 Implementation Details.
@@
4.2 Evaluation of Mapping and Tracking
4.2.1 Evaluation on Replica [23].
We compare the reconstruction performance on Replica [23] only with Co-SLAM [28] and ESLAM [12] as they significantly outperform previous methods [25, 35, 33]. For quantitative analysis, we run each method five times and report the average results. As shown in Tab. 1, our approach outperforms Co-SLAM [28] on all scenes and shows competitive performance with ESLAM [12]. Due to the use of a joint coordinate and parametric encoding, Co-SLAM [28] tends to produce over-smoothed surfaces, which leads to the amplification of reconstruction error. Although ESLAM [12] achieves high overall accuracy, it falls short in preserving surface details. To further highlight our method’s superiority in capturing scene details, we compare the $Completion\ ratio\ [<1cm\%]$ with ESLAM [12] in Tab. 2. Qualitative results in Fig. 3 also demonstrate that MUTE-SLAM effectively reconstructs detailed environmental geometry with fewer artifacts.
Table 1: Quantitative results of reconstruction on Replica [23] dataset. The best results are bolded. Our proposed MUTE-SLAM outperforms Co-SLAM [28] and shows better or on-par performance compared to ESLAM [12].
Metrics Method Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4 Avg.
Depth L1 (cm) $\downarrow$ ESLAM 0.95 0.98 1.37 0.81 1.27 1.60 1.49 1.06 1.19
Co-SLAM 1.47 1.44 3.70 1.99 1.71 7.76 5.42 2.73 3.22
Ours 0.93 0.95 1.35 0.78 1.31 1.56 1.69 0.88 1.18
Acc. (cm) $\downarrow$ ESLAM 1.04 0.80 0.90 0.79 0.83 0.96 1.16 1.06 0.94
Co-SLAM 1.04 0.91 1.01 0.88 0.70 1.86 1.94 1.12 1.18
Ours 1.03 0.81 0.90 0.78 0.68 0.93 1.13 1.04 0.91
Comp. (cm) $\downarrow$ ESLAM 1.03 0.83 0.99 0.82 0.72 0.98 1.16 1.12 0.96
Co-SLAM 1.06 0.93 1.21 0.93 0.77 1.34 1.46 1.26 1.12
Ours 1.01 0.81 0.98 0.80 0.73 0.97 1.18 1.08 0.95
Comp. Ratio (%) $\uparrow$ ESLAM 99.71 99.70 99.08 99.46 99.40 99.35 98.98 98.87 99.31
Co-SLAM 99.45 99.65 96.99 98.79 99.20 98.53 97.68 97.59 98.49
Ours 99.77 99.79 99.22 99.51 99.38 99.22 98.84 99.04 99.34
Table 2: Comparison of $Completion\ ratio\ [<1cm\%]$ with ESLAM [12]. Our method demonstrates superior performance in all scenes, indicative of our approach’s ability to accurately preserve details and minimize artifacts in reconstructed scene structures.
Metrics Method Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4 Avg.
Comp. Ratio $\uparrow$ ESLAM 55.21 72.77 65.94 76.30 86.17 62.90 49.05 54.55 65.36
@@ 4.2 Evaluation of Mapping and Tracking
Figure 3: Qualitative reconstruction results on Replica [23]. Our reconstructed mesh achieves a high level of completion and captures finer details of the environment, which are highlighted with green colored boxes across all methods.
4.2.2 Evaluation on ScanNet [6].
Table 3: Quantitative results of ATE RMSE (cm) on ScanNet [6]. While ESLAM [12] yields the best overall performance, our method surpasses it in several scenes and secures the second-best overall result.
@@
NICE-SLAM [35] 8.61 12.24 8.04 10.27 13.01 5.55 9.62
Vox-Fusion [33] 8.42 9.20 7.42 6.60 12.13 5.51 8.21
ESLAM [12] 7.37 8.31 7.59 6.45 9.29 5.65 7.44
Co-SLAM[28] 7.77 12.52 9.39 6.34 12.35 7.65 9.34
@@
We assessed camera tracking accuracy on six real-world scenes from the ScanNet dataset [6] and report the average ATE RMSE [24] across five runs for each scene and method in Tab. 3. Our approach, MUTE-SLAM, exhibits competitive results in these tests. Notably, even without pre-defined scene boundaries, MUTE-SLAM consistently outperforms Co-SLAM [28] in all tested scenes. While ESLAM [12] achieves the best overall performance, it does so with twice the average processing time as ours (Tab. 6). Due to the incomplete nature of ScanNet [6] meshes, we present only qualitative reconstruction results in Fig. 4. These results highlight MUTE-SLAM’s ability to capture finer details and achieve a high level of completeness in reconstructions.
4.2.3 Evaluation on TUM RGB-D [24].
Table 4: Quantitative results of ATE RMSE (cm) on TUM RGB-D [24] dataset. Instance of failure is denoted as N/A. Our proposed MUTE-SLAM outperforms NICE-SLAM [35] and Co-SLAM [28], while demonstrating competitive performance and superior robustness compared to ESLAM [12].
@@
NICE-SLAM [35] 2.7 1.8 3.0
ESLAM [12] 2.5 N/A 2.8
Co-SLAM[28] 2.9 1.8 2.9
@@
To further evaluate the tracking accuracy of MUTE-SLAM, we conducted experiments on real-world scenes from the TUM RGB-D dataset [24], with results averaged over five runs. Noted that ESLAM [12] runs unsuccessfully on the ’fr2/xyz’ scene and has been consequently excluded from these comparative results. As shown in Tab. 4, our quantitative analysis reveals that MUTE-SLAM not only outperforms NICE-SLAM [35] and Co-SLAM [28] but also demonstrates competitive performance and superior robustness compared to ESLAM [12], which takes hours to run on this dataset while ours only takes a few minutes.
4.2.4 Evaluation on Apartment [35] dataset.
To demonstrate the effectiveness of our method in large scale indoor scenarios, we evaluate the tracking and surface reconstruction performance on Apartment dataset provided by NICE-SLAM [35]. As illustrated in Tab. 5, our method yields reasonable tracking performance. It should be emphasized that our method runs the fastest on this dataset, as discussed in Sec. 4.3.
Table 5: Quantitative results of ATE RMSE (cm) on Apartment [35] dataset. The failed instance is denoted as N/A. Our approach generates reasonable tracking result on this dataset.
Method NICE. [35] Vox. [33] ESLAM [12] Co.[28] Ours

### v3

@@ Abstract
We introduce MUTE-SLAM, a real-time neural RGB-D SLAM system employing multiple tri-plane hash-encodings for efficient scene representation. MUTE-SLAM effectively tracks camera positions and incrementally builds a scalable multi-map representation for both small and large indoor environments. As previous methods often require pre-defined scene boundaries, MUTE-SLAM dynamically allocates sub-maps for newly observed local regions, enabling constraint-free mapping without prior scene information. Unlike traditional grid-based methods, we use three orthogonal axis-aligned planes for hash-encoding scene properties, significantly reducing hash collisions and the number of trainable parameters. This hybrid approach not only ensures real-time performance but also enhances the fidelity of surface reconstruction. Furthermore, our optimization strategy concurrently optimizes all sub-maps intersecting with the current camera frustum, ensuring global consistency. Extensive testing on both real-world and synthetic datasets has shown that MUTE-SLAM delivers state-of-the-art surface reconstruction quality and competitive tracking performance across diverse indoor settings. The code is available at https://github.com/lumennYan/MUTE_SLAM.
Index Terms:
neural RGB-D SLAM, tri-plane hash-encoding, multi-map representation
I Introduction
@@
With the advent of Neural Radiance Field (NeRF) [13], efforts have been made to integrate it into SLAM systems due to its capability to render novel views [14, 15, 16, 17] and reconstruct 3D surfaces [18, 19, 20, 21]. NeRF-based SLAM methods iMAP [22] and NICE-SLAM [23], have demonstrated their applicability across various scenes and ability to predict the appearance of unobserved areas, although their computational demands hinder real-time application. Recent works [24, 25] utilize hash-encoded voxel grids [26] to accelerate convergence and enhance detail fidelity. Despite hash collisions can be implicitly mitigated by the original design of Instant-NGP [26], the reconstructed mesh still suffers from aliasing. As proved in [27, 28, 29, 30, 29], projecting spatial points into a tri-plane reprensentation can efficiently reduce scene parameters while preserving geometric and color information. Therefore following [30], we leverage a tri-plane hash-encoding method to store scene features and minimize hash collisions. Moreover, current Nerf-based SLAM systems [22, 23, 28, 24, 31] require pre-set scene boundaries, limiting their application in unknown environments. Some [25, 32] address this with octree-based voxel grids, but they still necessitate an initially defined loose boundary and struggle to reconstruct beyond these limits. Although neural point cloud-based method [33] does not have such a concern, the point cloud-based representation requires large time and memory consumption, making it impractical for real-time usage. Our proposed MUTE-SLAM overcomes this by introducing a multi-map-based scene representation. Through dynamically allocating new sub-maps upon detecting new areas, MUTE-SLAM can be deployed in environments of any size, given reasonable RGB-D sensor observations, while maintaining reasonable runtime and memory overhead.
@@
II Related Works
II-A Dense Visual SLAM
@@
II-B Neural Implicit SLAM
@@
Recent advancements in NeRF-based SLAM have improved tracking, mapping, and running speed. ESLAM [28] stores features on multi-scale axis-aligned planes and employs rendering based on TSDF. Co-SLAM [24] combines coordinate and hash-encodings for input points to achieve both smooth meshes and fast convergence, and introduces a real-time global bundle adjustment mechanism using a ray list sampled from past keyframes. GO-SLAM [31] and H2-mapping [25] combine traditional SLAM modules with neural mapping, achieving high tracking accuracy. Despite these improvements, scalability remains a challenge. Our work addresses this by introducing a multi-map solution with an accompanying optimization strategy, which requires no pre-set boundaries. Besides, we represent each sub-map with a tri-plane hash-encoding which allows for fast convergence and detailed surface reconstruction.
@@
III Method
The overview of MUTE-SLAM is illustrated in fig. 2. Given an input RGB-D stream $\{I_{i},D_{i}\}_{i=1}^{M}$, the system tracks their 6-DOF camera poses $\{T_{i}:R_{i}|t_{i}\}_{i=1}^{M}$ and conducts mapping to optimize a multi-map implicit scene representation $\{E_{k}\}_{k=1}^{S}$. When the camera captures a new region, a corresponding sub-map is created to represent it, as detailed in section III-A. Each local map $E_{k}$ encodes a point coordinate $\bm{p}$ within its domin with three orthogonal TSDF feature planes and three color planes, as described in section III-B. These features from the sub-maps are decoded into TSDF and color using two separate MLPs, initiating the volume rendering process (section III-C). The rendered depth and color images are subsequently utilized to jointly optimize the sub-maps and camera poses. Additionally, periodic global bundle adjustments are implemented to ensure global consistency(section III-D).
III-A Multi-map Scene Representation
@@
III-B Tri-plane Hash Encoding
@@
The TSDF and color values are then utilized in the volume rendering module. The plane-based representation, growing quadratically with scene size, results in fewer hash table queries and hence fewer collisions compared to grid representations, given equal hash table sizes. Furthermore, since a point’s feature vector is a composite of inputs from three distinct encoders, the probability of encountering conflicts across all encoders is significantly reduced. In situations where collisions do occur in one encoder, the impact is mitigated by the inputs from the other encoders, thus lessening the overall adverse effect. section IV-D demonstrates the effectiveness of our tri-plane hash-encoding.
III-C TSDF-based Volume Rendering
@@
For $\omega_{n}$, as in [19], it is derived from the volume density $\sigma(\bm{p_{n}})$ and TSDF value $s(\bm{p_{n}})$:
@@
III-D Tracking and Mapping
III-D1 Loss Functions.
We apply four loss functions to optimize the scene representation, MLPs and camera poses: RGB loss, depth loss, TSDF loss and free-space loss. Once a batch of rays $R$ are selected, the RGB and depth loss are obtained as $l_{2}$ errors between rendered and ground truth values:
@@ III-D Tracking and Mapping
For points near the surface ($|d_{p}-d(r)|<=\delta$), similar to [28], we further split them into two parts to obtain the TSDF loss:
@@
IV Experiments.
IV-A Experimental Setup.
IV-A1 Baselines
We choose state-of-art NeRF-based dense SLAM approaches ESLAM [28], Co-SLAM [24] and Point-SLAM [33] as our main baselines for both surface reconstruction and camera tracking. To better evaluate our proposed MUTE-SLAM on pose estimation, we also compare with previous methods NICE-SLAM [23] and Vox-Fusion [32]. We run these methods using the default settings provided in their open-source code. We do not compare with[25] and [31] as they both use traditional SLAM modules for camera poses estimation.
IV-A2 Datasets
@@
IV-A3 Metrics
For surface reconstruction, We adopt four evaluation metrics: $Depth\ L1\ [cm]$, $Accuracy\ [cm]$, $Completion\ [cm]$ and $Completion\ ratio\ [<5cm\%]$. Additionally, to underscore the ability of our method to produce detailed geometry compared to ESLAM [28], we also incorporate the $Completion\ ratio\ [<1cm\%]$ metric. Following [18, 37, 28], before evaluation, we remove faces that are not inside any camera frustum or are occluded in all RGB-D frames from the reconstructed mesh. For the evaluation of camera tracking, we employ $ATE\ RMSE\ [cm]$ [34].
IV-A4 Implementation Details
@@
IV-B Evaluation of Mapping and Tracking
IV-B1 Evaluation on Replica [35]
We compare the reconstruction performance on Replica [35] only with Co-SLAM [24], Point-SLAM [33] and ESLAM [28] as they significantly outperform previous methods [22, 23, 32]. However, the origin setting of Point-SLAM takes hours on Replica, which is unfair for other baselines as they only require a few minutes. Thus we modified the number of rays to 2000 and iterations to 10 for Point-SLAM, and denote the modified version as $\text{Point}^{\ast}$. Note that $\text{Point}^{\ast}$ [sandstrom2023poin] still takes two times longer than other methods. For quantitative analysis, we run each method five times and report the average results. As shown in table I, our approach outperforms Co-SLAM [24] and $\text{Point}^{\ast}$ [33] on all scenes and shows competitive performance with ESLAM [28]. Due to the use of a joint coordinate and parametric encoding, Co-SLAM [24] tends to produce over-smoothed surfaces, which leads to the amplification of reconstruction error. Point-SLAM [33] can reconstruct fine scene contents as in the origin setting, but is impractical in real-time use. Although ESLAM [28] achieves high overall accuracy, it falls short in preserving surface details. To further highlight our method’s superiority in capturing scene details, we compare the $Completion\ ratio\ [<1cm\%]$ with ESLAM [28] in table II. Qualitative results in fig. 3 also demonstrate that MUTE-SLAM effectively reconstructs detailed environmental geometry with fewer artifacts. Because the modified version $\text{Point}^{\ast}$ [sandstrom2023poin] performs poorly, we do not show its qualitative results.
Figure 3: Qualitative reconstruction results on Replica.
TABLE I: Quantitative results of reconstruction on Replica [35] dataset.
Method Depth L1 (cm) $\downarrow$ Acc. (cm) $\downarrow$ Comp. (cm) $\downarrow$ Cp. Ratio (%) $\uparrow$
ESLAM 1.19 0.94 0.96 99.31
Co-SLAM 3.22 1.18 1.12 98.49
$\text{Point}^{\ast}$ 8.39 3.72 2.10 94.09
Ours 1.18 0.91 0.95 99.34
TABLE II: Comparison of $Completion\ ratio\ [<1cm\%]$ with ESLAM[28].
Method Room0 Room1 Room2 Office0 Office1 Office2 Office3 Office4 Avg.
ESLAM 55.21 72.77 65.94 76.30 86.17 62.90 49.05 54.55 65.36
@@ IV-B1 Evaluation on Replica [35]
IV-B2 Evaluation on ScanNet[36]
TABLE III: Quantitative results of ATE RMSE (cm) on ScanNet [36].
@@ IV-B2 Evaluation on ScanNet[36]
NICE-SLAM 8.61 12.24 8.04 10.27 13.01 5.55 9.62
Vox-Fusion 8.42 9.20 7.42 6.60 12.13 5.51 8.21
ESLAM 7.37 8.31 7.59 6.45 9.29 5.65 7.44
Co-SLAM 7.77 12.52 9.39 6.34 12.35 7.65 9.34
Point-SLAM 10.24 7.81 8.65 22.16 14.77 9.54 12.19
@@
We assessed camera tracking accuracy on six real-world scenes from the ScanNet dataset [36] and report the average ATE RMSE [34] across five runs for each scene and method in table III. Our approach, MUTE-SLAM, exhibits competitive results in these tests. Notably, even without pre-defined scene boundaries, MUTE-SLAM consistently outperforms Co-SLAM [24] in all tested scenes. While ESLAM [28] achieves the best overall performance, it does so with twice the average processing time as ours (table VI). Our approach not only surpasses ESLAM in several scenes but also secures the second-best overall result.Due to the incomplete nature of ScanNet [36] meshes, we present only qualitative reconstruction results in fig. 4. These results highlight MUTE-SLAM’s ability to capture finer details and achieve a high level of completeness in reconstructions.
IV-B3 Evaluation on TUM RGB-D [34]
TABLE IV: Quantitative results of ATE RMSE (cm) on TUM RGB-D[34].
@@ IV-B3 Evaluation on TUM RGB-D [34]
NICE-SLAM 2.7 1.8 3.0
ESLAM 2.5 N/A 2.8
Co-SLAM 2.9 1.8 2.9
@@
BAD-SLAM 1.7 1.1 1.7
Kintinuous 3.7 2.9 3.0
ORB-SLAM2 1.6 0.4 1.0
To further evaluate the tracking accuracy of MUTE-SLAM, we conducted experiments on real-world scenes from the TUM RGB-D dataset [34], with results averaged over five runs. Instance of failure is denoted as ‘N/A’. Noted that ESLAM [28] runs unsuccessfully on the ’fr2/xyz’ scene and has been consequently excluded from these comparative results. As shown in table IV, our quantitative analysis reveals that MUTE-SLAM not only outperforms NICE-SLAM [23] and Co-SLAM [24] but also demonstrates competitive performance and superior robustness compared to ESLAM [28], which takes hours to run on this dataset while ours only takes a few minutes. We also compare with some traditional SLAM methods [9, 5, 38] on this dataset. While NeRF-based SLAM methods still lag behind them, MUTE-SLAM narrows the gap.
IV-B4 Evaluation on Apartment [23]
To demonstrate the effectiveness of our method in large scale indoor scenarios, we evaluate the tracking and surface reconstruction performance on Apartment dataset provided by NICE-SLAM [23]. The failed instance is denoted as N/A. As illustrated in table V, our method yields reasonable tracking performance. It should be emphasized that our method runs the fastest on this dataset, as discussed in section IV-C.
TABLE V: Quantitative results of ATE RMSE (cm) on Apartment [23] dataset.
Method NICE. Vox. ESLAM Co. Ours